nolegirl24:

Choose the correct simplification of the expression (3xy^4)^2(y^2)^3. 6x^2y^14 9x^2y^14 9x^3y^11 6x^3y^11

dude:

\((3xy^4)^2(y^2)^3\) When raising a power to a power, multiply the exponents

nolegirl24:

Okay writing this down sorry. I dont retain information that good

dude:

Ah okay

nolegirl24:

So it would have to be like this? 4*2*2*3

dude:

\(\large (3xy^4)^2=3^2~x^{1\times2}~y^{4\times 2}\\ (y^2)^3=y^{2\times 3}\)

nolegirl24:

Okay wrote it down

nolegirl24:

So would it end up being \[9x^2y^8\] and \[y^6\]

dude:

Yes Now add the \(y^6\) and the \(y^8\)

nolegirl24:

y^14?

dude:

Yes

nolegirl24:

\[9x^2y^14\]

dude:

Yes

nolegirl24:

Thankssssss